# A Prospect-Theoretic Policy Gradient Framework for Behaviorally Nuanced Reinforcement Learning

### v3

@@ Abstract
Classical reinforcement learning (RL) typically assumes rational decision-making based on expected utility theory. However, this model has been shown to be empirically inconsistent with actual human preferences, as evidenced in psychology and behavioral economics. Cumulative Prospect Theory (CPT) provides a more nuanced model for human-based decision-making, capturing diverse attitudes and perceptions toward risk, gains, and losses. While prior work has integrated CPT with RL to solve CPT policy optimization problems, the understanding and impact of this formulation remain limited. Our contributions are as follows: (a) we derive a novel policy gradient theorem for CPT objectives, generalizing the foundational result in standard RL, (b) we design a model-free policy gradient algorithm for solving the CPT-RL problem, (c) we analyze our policy gradient estimator and prove asymptotic convergence of the algorithm to first-order stationary points, and (d) test its performance through simulations. Notably, our first-order policy gradient algorithm scales better than existing zeroth-order methods to larger state spaces. Our theoretical framework offers more flexibility to advance the integration of behavioral decision-making into RL.
@@ 1 Introduction
While risk-sensitive RL extends beyond risk-neutral settings to capture individual risk preferences (using e.g. variance or conditional value at risk metrics), it remains limited in fully representing the complexities of human decision-making and fail to capture the psychological nuances of how humans perceive risks, gains, and losses, often leading to decisions misaligned with human preferences. Specifically, risk-sensitive RL fails to account for the asymmetric perception of gains and losses, as well as the probability distortions inherent in human cognition, such as the tendency to overestimate rare events and underestimate frequent ones (see App. C for an illustration in a simulation). These psychological and cognitive biases, critical to understanding real-world human decisions under uncertainty, are beyond the scope of traditional risk-sensitive RL frameworks, necessitating a more comprehensive approach.
@@
Cumulative Prospect Theory (CPT), developed by the pioneering behavioral psychologist Daniel Kahneman111was awarded the Nobel Prize in Economic Sciences in 2002 ‘for having integrated insights from psychological research into economic science, especially concerning human judgment and decision-making under uncertainty’. together with his colleague Amos Tversky in their seminal works combining cognitive psychology and economics (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992), provides a more nuanced understanding of human decision-making to explain several empirical observations invalidating the standard expected utility theory. CPT models how individuals perceive outcomes asymmetrically, being risk-averse in the domain of gains and risk-seeking in the domain of losses, and distort probabilities to reflect cognitive biases. These insights have led to widespread applications of CPT in stateless static settings in high-stakes domains such as healthcare in psychiatry Sip et al. (2018); George et al. (2019); Mkrtchian et al. (2023), chronic diseases treatment Zhao et al. (2023) and emergency decision making Sun et al. (2022) where taking into account psychological factors is of paramount importance, as well as other application domains such as energy Ebrahimigharehbaghi et al. (2022); Dorahaki et al. (2022) and finance Ladrón de Guevara Cortés et al. (2023); Luxenberg et al. (2024) to name a few. However, these applications often overlook the sequential decision-making nature of many real-world problems, where the outcome of one decision can affect future choices, a critical aspect of RL.
@@
The integration of CPT into RL provides a promising avenue for behaviorally-aligned sequential decision-making, as it allows RL agents to consider both risk preferences and probability distortions in dynamic environments. While a few isolated recent works have explored this integration (L.A. et al., 2016; Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Danis et al., 2023; Foo et al., 2023), the understanding and practical impact of CPT in RL remains limited. Specifically, the computational challenges and theoretical underpinnings of CPT-based RL models are still underexplored, and the practical effectiveness of such models in sequential decision-making problems is not well understood, especially due to the lack of efficient policy gradient algorithms like in standard RL.
@@
(ii) Policy gradient algorithm for CPT-RL. Building on our policy gradient theorem, we design a policy gradient (PG) algorithm to solve the CPT policy optimization problem. To the best of our knowledge, this is the first PG algorithm for CPT-RL using first-order information which does not rely on zeroth-order policy gradient estimation.
(iii) Convergence and sample complexity. We show that our PG estimator is consistent and we analyze its sample complexity. Notably, our sample complexity only scales logarithmically with the dimension of the policy parameter, in contrast to existing zeroth-order PG estimators. We further show that the iterates of our PG algorithm converge asymptotically to the set of stationary points of the CPT-RL objective which is highly non-convex. In particular, unlike prior work, we do not require increasing batch sizes for our PG estimator to guarantee convergence.
(iv) Simulations. We test our PG algorithm on several applications to illustrate the flexibility of CPT-RL compared to standard and risk-sensitive RL in leading to more nuanced behavior. We also compare the performance of our PG algorithm to the previously proposed zeroth order algorithm to show the robustness of our algorithm to larger MDPs.
This paper contributes to the development of more efficient algorithms for human-centered sequential decision-making with models able to incorporate cognitive and psychological biases.
2 From Classical to CPT-RL
2.1 MDPs, CPT Values and Notation
@@
$\Pi_{M,S}\subseteq\Pi_{M,NS}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}$ and $\Pi_{M,S}\subseteq\Pi_{\Sigma,S}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}\,$, see Fig. 4 in App. B.
@@
(c) A probability distortion function $w:[0,1]\to[0,1]$. This is a continuous non-decreasing weight function that distorts the probability distributions of the gain and loss variables. This function typically captures the human tendency to overestimate the probability of rare events and underestimate the probability of more certain ones. Similarly to the utility function, we denote by $w^{+}$ (resp. $w^{-}$) the function that warps the cumulative distribution function for gains (resp.for losses). Both functions are required to be defined on $[0,1]$, with values in $[0,1]$ and to be non-decreasing, continuous, with $w^{+}(0)=w^{-}(0)=0$ and $w^{-}(1)=w^{-}(1)=1$. Examples of such weights functions in the literature include $w:p\mapsto p^{\eta}(p^{\eta}+(1-p)^{\eta})^{-\frac{1}{\eta}}$ (Kahneman and Tversky, 1979) and $w:p\mapsto\exp(-(-\ln p)^{\eta})$ (Prelec, 1998) where $\eta\in(0,1)$ is a hyperparameter. We refer the reader to App. G.3 for examples and plots of utility and probability weight functions.
@@
$\mathbb{C}(X)=\int_{0}^{+\infty}w^{+}(\mathbb{P}(u^{+}(X)>z))dz-\int_{0}^{+\infty}w^{-}(\mathbb{P}(u^{-}(X)>z))dz\,$, (1)
where appropriate integrability assumptions are assumed. While the CPT value $\mathbb{C}(X)$ accounts for the human agent’s distortions in perception, it also recovers the expectation $\mathbb{E}(X)$ with weight functions $w^{+},w^{-}$ and utility functions $u^{+}$ (resp. $u^{-}$) restricted to $\mathbb{R}_{+}$ (resp. $\mathbb{R}_{-}$) are set to be the identity functions. In addition, several risk measures such as variance, Conditional Value at Risk (CVaR) and distortion risk measures are also particular cases of CPT values with discontinuous probability weighting functions (see App. G and Table 1 therein).
@@ 2.2 CPT-RL Problem Formulation
For an illustration of the CPT-RL problem formulation and its elements in a personal treatment for pain management application, see App. E.5. In the particular case of CPT-PO in which $w^{+},w^{-}$ are set to the identity, namely the Expected Utility Policy Optimization (EUT-PO) objective, only returns are distorted by the utility function:
@@
Scope of this work. We assume the utility and weighting functions are known, e.g., from domain experts, surveys, prior empirical studies on target user groups, or behavioral studies (Mkrtchian et al., 2023; George et al., 2019; Sip et al., 2018), see App. G.7 for more details on the choice of the utility function. Our goal is to align the agent’s behavior with the given preferences (encoded in utility and weighting functions) by optimizing for the CPT value of returns.
We leave the question of discovering or inferring human preferences (e.g. RL from Human Feedback (RLHF)) to future work. See section F for an extended discussion comparing CPT-RL to RLHF as preference learning paradigms, their pros and cons and opportunities for future work in combining them as they are not mutually exclusive. See also App. E for applications.
@@ 2.3 Peculiarities of Optimal Policies in CPT-RL
In this section, we highlight the properties of optimal policies to CPT-PO when they exist and contrast them with classical RL and EUT-PO (details in App. H):
The need for stochastic policies. In MDPs, optimal deterministic stationary policies always exist; in CPT-RL, optimal policies may need to be stochastic. In general, stochasticity of the policy is essential in solving CPT-PO. See App. H.1 for a proof.
@@ 3.1 CPT Policy Gradient Theorem
where $R(\tau):=\sum_{t=0}^{H-1}r_{t}\,$ with $\tau:=(s_{t},a_{t},r_{t})_{0\leq t\leq H-1}$ is a trajectory of length $H$ generated from the MDP by following policy $\pi_{\theta}\,$ [^3] The integral $\varphi(R(\tau))$ is finite under our continuity assumptions since the return $R(\tau)$ is bounded. and for all $v\in\mathbb{R},\varphi(v):=\int_{z=0}^{\max(v,0)}w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z))dz$ $\int_{z=0}^{\max(-v,0)}w_{-}^{\prime}(\mathbb{P}(u^{-}((R(\tau)>z))dz,$ and $w_{+}^{\prime},w_{-}^{\prime}$ denoting the derivatives.
We provide a few comments regarding this result. Theorem 3 recovers the celebrated policy gradient theorem for standard RL (Sutton et al., 1999) by setting $w_{+}$ (resp. $w_{-}$) to the identity function (on $\mathbb{R}_{+}$ (resp. $\mathbb{R}^{-}$) in which case $w_{+}^{\prime}$ is the constant function equal to $1$ and hence $\varphi(R(\tau))=R(\tau)\,.$ We stated the theorem in the general setting where the policy is non-Markovian. In practice, it is also possible to use a parametrization of a smaller policy set such as $\Pi_{\Sigma,NS}$ or even $\Pi_{M,S}$ in which the policy is a function of $(t,s_{t},\sum_{k=0}^{t-1}r_{k})$ or only $s_{t}$ respectively.
@@ 3.2 Stochastic PG Algorithm for CPT-RL
7:  Compute and order $R(\tau_{j})$, label them as   $R(\tau_{[1]})<R(\tau_{[2]}<\cdots<R(\tau_{[n]})$
@@ 4 Asymptotic Convergence and Sample Complexity
Let the assumptions of Prop. 5 hold. Suppose in addition that the score function is bounded, i.e. there exists $M_{\psi}>0$ s.t. $\|\nabla\ln\pi_{\theta}(a|s)\|_{2}\leq M_{\psi}$ for all $\theta\in\mathbb{R}^{d},(s,a)\in\mathcal{S}\times\mathcal{A}$. Then there exists an absolute constant $c>0$ s.t. for any $\varepsilon>0,\delta\in(0,1)$ and for all $\theta\in\mathbb{R}^{d},$ we have $\|\hat{\nabla}_{n,m}J(\theta)-\nabla J(\theta)\|_{2}\leq\varepsilon$ with probability at least $1-\delta$, if $n\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(1/\delta)}{\varepsilon^{2}}$ and $m\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(2d/\delta)}{\varepsilon^{2}}\,.$
@@
We close this section by discussing the convergence of the policy parameter sequence produced by Algorithm 1. The CPT-RL objective is non-convex in the policy parameter due to non-convexity of both utility and probability weighting functions. While this lack of structure makes global optimality out of reach, we can still obtain a standard asymptotic convergence result towards the set of stationary points using the machinery of stochastic approximation (see e.g. (Borkar and Borkar, 2008; Benaïm, 2006)).

### v4

# Policy Gradients for Cumulative Prospect Theory in Reinforcement Learning

@@ Abstract
We derive a policy gradient theorem for Cumulative Prospect Theory (CPT) objectives in finite-horizon Reinforcement Learning (RL), generalizing the standard policy gradient theorem and encompassing distortion-based risk objectives as special cases. Motivated by behavioral economics, CPT combines an asymmetric utility transformation around a reference point with probability distortion. Building on our theorem, we design a first-order policy gradient algorithm for CPT-RL using a Monte Carlo gradient estimator based on order statistics. We establish statistical guarantees for the estimator and prove asymptotic convergence of the resulting algorithm to first-order stationary points of the (generally non-convex) CPT objective. Simulations illustrate qualitative behaviors induced by CPT and compare our first-order approach to existing zeroth-order methods.
@@ 1 Introduction
While risk-sensitive RL extends beyond risk-neutral settings to capture individual risk preferences (using e.g. variance or conditional value at risk metrics), many commonly used risk-sensitive objectives capture particular aspects of risk preferences (e.g., tail risk) but do not simultaneously model gain–loss asymmetry around a reference point together with probability distortion. In particular, many standard risk-sensitive criteria do not model the asymmetric perception of gains and losses, as well as the probability distortions inherent in human cognition, such as the tendency to overestimate rare events and underestimate frequent ones (see App. F for an illustration in a simulation). These behavioral phenomena, relevant in human-facing decision problems under uncertainty, are beyond the scope of traditional risk-sensitive RL frameworks, necessitating a more comprehensive approach.
@@
Cumulative Prospect Theory (CPT), introduced by Kahneman and Tversky in their seminal works combining cognitive psychology and economics (Kahneman and Tversky, 1979; Tversky and Kahneman, 1992), provides a descriptive behavioral model of decision-making under risk to explain several empirical observations that challenge the standard expected utility theory. CPT models how individuals perceive outcomes asymmetrically, being risk-averse in the domain of gains and risk-seeking in the domain of losses, and distort probabilities to reflect cognitive biases. These insights have led to widespread applications of CPT in stateless static settings in domains such as healthcare in psychiatry (Sip et al., 2018; George et al., 2019; Mkrtchian et al., 2023), chronic diseases treatment (Zhao et al., 2023) and emergency decision making (Sun et al., 2022) where modeling behavioral factors can be important, as well as other application domains such as energy (Ebrahimigharehbaghi et al., 2022; Dorahaki et al., 2022) and finance (Ladrón de Guevara Cortés et al., 2023; Luxenberg et al., 2024) to name a few. However, these applications often overlook the sequential decision-making nature of many real-world problems, where the outcome of one decision can affect future choices, a critical aspect of RL.
@@
The integration of CPT into RL provides a promising avenue for behaviorally-aligned sequential decision-making, as it allows RL agents to consider both risk preferences and probability distortions in dynamic environments. While a few isolated recent works have explored this integration (L.A. et al., 2016; Borkar and Chandak, 2021; Ramasubramanian et al., 2021; Danis et al., 2023; Foo et al., 2023), the understanding and practical impact of CPT in RL remains limited. Specifically, the computational challenges and theoretical underpinnings of CPT-based RL models are still underexplored.
@@
(ii) Policy gradient algorithm for CPT-RL. Building on our policy gradient theorem, we design a policy gradient (PG) algorithm to solve the CPT policy optimization problem. Our PG algorithm for CPT-RL uses first-order information and does not rely on zeroth-order policy gradient estimation.
(iii) Convergence and sample complexity. We show that our PG estimator is consistent and we analyze its sample complexity. Notably, our sample complexity scales logarithmically in the policy-parameter dimension (under our estimator and regularity assumptions), in contrast to existing zeroth-order PG estimators. We further show that the iterates of our PG algorithm converge asymptotically to the set of stationary points of the CPT-RL objective which is typically non-convex in the policy parameters. In particular, our analysis does not require increasing batch sizes to establish asymptotic convergence.
(iv) Simulations. We test our PG algorithm on several applications to illustrate the flexibility of CPT-RL compared to standard and risk-sensitive RL in leading to more nuanced behavior. We also compare the performance of our PG algorithm to the previously proposed zeroth order algorithm as the problem size increases in our simulation settings.
Closest related work. Closest to our work are policy gradient methods for distortion risk measures (DRMs) (Vijayan and L.A, 2024), which optimize distorted-expectation objectives via first-order gradients. CPT shares the probability distortion component but additionally incorporates a reference point and an asymmetric utility transformation (and potentially separate gain/loss distortions), enabling objectives beyond distortion-only criteria. Our policy gradient expression reduces to DRM policy gradients under the corresponding restrictions (e.g., identity utility and no reference point), while providing a first-order method for the more general CPT objective. We provide additional discussion after Theorem 3. A more comprehensive related work discussion can be found in section 6 and App. B.
2 From Classical RL To CPT-RL
2.1 MDPs, CPT and Notation
@@
$\Pi_{M,S}\subseteq\Pi_{M,NS}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}$ and $\Pi_{M,S}\subseteq\Pi_{\Sigma,S}\subseteq\Pi_{\Sigma,NS}\subseteq\Pi_{NM}\,$, see Fig. 4 in App. A.
@@
(c) A probability distortion function $w:[0,1]\to[0,1]$. This is a continuous non-decreasing weight function that distorts the probability distributions of the gain and loss variables. Monotonicity here preserves ordering: it is natural to suppose that events with (truly) higher probability remain perceived as more probable than less truly probable ones after distortion. This distortion function $w$ typically captures the human tendency to overestimate the probability of rare events and underestimate the probability of more certain ones. Similarly to the utility function, we denote by $w_{+}$ (resp. $w_{-}$) the function that warps the cumulative distribution function for gains (resp.for losses). Both functions are required to be defined on $[0,1]$, with values in $[0,1]$ and to be non-decreasing, continuous, with $w_{+}(0)=w_{-}(0)=0$ and $w_{+}(1)=w_{-}(1)=1$. Examples of such weights functions in the literature include $w:p\mapsto p^{\eta}(p^{\eta}+(1-p)^{\eta})^{-\frac{1}{\eta}}$ (Kahneman and Tversky, 1979) and $w:p\mapsto\exp(-(-\ln p)^{\eta})$ (Prelec, 1998) where $\eta\in(0,1)$ is a hyperparameter. We refer the reader to App. I.3 for examples and plots of utility and probability weight functions.
@@
$\mathbb{C}(X)=\int_{0}^{+\infty}w_{+}(\mathbb{P}(u^{+}(X)>z))dz-\int_{0}^{+\infty}w_{-}(\mathbb{P}(u^{-}(X)>z))dz\,$, (1)
where appropriate integrability assumptions are assumed. While the CPT value $\mathbb{C}(X)$ accounts for the human agent’s distortions in perception, it also recovers the expectation $\mathbb{E}(X)$ with weight functions $w_{+},w_{-}$ and utility functions $u^{+}$ (resp. $u^{-}$) restricted to $\mathbb{R}_{+}$ (resp. $\mathbb{R}_{-}$) are set to be the identity functions. In addition, several risk measures such as variance, Conditional Value at Risk (CVaR) and distortion risk measures are also particular cases of CPT values with discontinuous probability weighting functions (see App. I and Table 1 therein).
@@ 2.2 CPT-RL Problem Formulation
For an illustration of the CPT-RL problem formulation and its elements in a personal treatment for pain management application, see App. G.5. In the particular case of CPT-PO in which $w_{+},w_{-}$ are set to the identity, namely the Expected Utility Policy Optimization (EUT-PO) objective, only returns are distorted by the utility function:
@@
Scope of this work. We assume that the utility and weighting functions are known, e.g., from domain experts, surveys, prior empirical studies on target user groups, or behavioral studies (Mkrtchian et al., 2023; George et al., 2019; Sip et al., 2018), see App. I.7 for more details on the choice of the utility function. Our goal is to align the agent’s behavior with the given preferences (encoded in utility and weighting functions) by optimizing for the CPT value of returns. Typically, utility and weight functions are chosen as the KT model and hyperparameters of this model are estimated from data. We leave the question of discovering or inferring human preferences (e.g. RL from Human Feedback (RLHF)) to future work. [^2] See section H for an extended discussion comparing CPT-RL to RLHF as preference learning paradigms, their pros and cons and opportunities for future work in combining them as they are not mutually exclusive. See also App. G for applications. Note that the transition model $\mathcal{P}$ and the reward function $r$ are still supposed to be unknown in our setting. In particular, we do not estimate them as in a model-based approach and our algorithm only uses sampled state-action-reward trajectories.
@@ 2.3 Peculiarities of Optimal Policies in CPT-RL
In this section, we highlight the properties of optimal policies to CPT-PO when they exist and contrast them with classical RL and EUT-PO (details in App. I.1):
The need for stochastic policies. In MDPs, optimal deterministic stationary policies always exist; in CPT-RL, optimal policies may need to be stochastic. In general, stochasticity of the policy is essential in solving CPT-PO. See App. C.1 for a proof.
@@ 3.1 CPT Policy Gradient Theorem
where $R(\tau):=\sum_{t=0}^{H-1}r_{t}\,$ with $\tau:=(s_{t},a_{t},r_{t})_{0\leq t\leq H-1}$ is a trajectory of length $H$ generated from the MDP by following policy $\pi_{\theta}\,$ [^3] The integral $\varphi(R(\tau))$ is finite under our continuity assumptions since the return $R(\tau)$ is bounded. and
$\varphi(R(\tau)):=\int_{z=0}^{u^{+}(R(\tau))}w_{+}^{\prime}(\mathbb{P}(u^{+}(R(\tau))>z))dz-\int_{z=0}^{u^{-}(R(\tau))}w_{-}^{\prime}(\mathbb{P}(u^{-}(R(\tau))>z))dz\,$, (2)
where $w_{+}^{\prime},w_{-}^{\prime}$ denote the derivatives of the functions $w_{+},w_{-}$ respectively.
We provide a few comments regarding this result:
  • Theorem 3 recovers the celebrated policy gradient theorem for standard RL (Sutton et al., 1999) by setting $w_{+}$ (resp. $w_{-}$) to the identity function (on $\mathbb{R}_{+}$ (resp. $\mathbb{R}^{-}$) in which case $w_{+}^{\prime}$ is the constant function equal to 11 and hence $\varphi(R(\tau))=R(\tau)\,.$
  • In the special case where (i) the CPT utilities are identity functions ($u^{+}(x)=x,u^{-}(x)=x$) and (ii) the gain/loss weighting functions derive from a single distortion function $g$ (i.e. $w_{+}(t)=1-g(1-t),w_{-}(t)=g(t)$), the CPT value reduces to a DRM objective (see App. I.4) and Theorem 3 recovers the DRM policy gradient of Vijayan and L.A (2024). CPT objectives are strictly more expressive than distortion risk metrics: DRMs impose the constraints $u^{+}(x)=x,u^{-}(x)=x,w_{+}(t)+w_{-}(1-t)=1$. This duality constraint implies that overweighting low probabilities of large losses implies a specific fixed way of transforming probabilities of large gains. As a consequence, behaviors like ‘extreme pessimism about rare losses and mild optimism about rare gains’ cannot be captured by DRMs. CPT is more expressive as it allows arbitrary $u_{+},u_{-},w_{+},w_{-}$ with no duality constraint. In particular, it captures reference dependence, loss aversion, unequal tail distortions and asymmetric risk attitudes. CPT uses different utility and probability weight functions for gains and loss regions (possibly asymmetric), allowing for more nuanced behaviors, e.g. risk seeking for gains and risk averse for losses at the same time.
  • We stated the theorem in the general setting where the policy is non-Markovian. In practice, it is also possible to use a parametrization of a smaller policy set such as $\Pi_{\Sigma,NS}$ or even $\Pi_{M,S}$ in which the policy is a function of $(t,s_{t},\sum_{k=0}^{t-1}r_{k})$ or only $s_{t}$ respectively.
@@ 3.2 Stochastic PG Algorithm for CPT-RL
7:  Compute and order $R(\tau_{j})$, label them as   $R(\tau_{[1]})<R(\tau_{[2]})<\cdots<R(\tau_{[n]})$
@@ 4 Asymptotic Convergence and Sample Complexity
Let the assumptions of Prop. 5 hold. Suppose in addition that the score function is bounded, i.e. there exists $M_{\psi}>0$ s.t. $\|\nabla\ln\pi_{\theta}(a|s)\|_{2}\leq M_{\psi}$ for all $\theta\in\mathbb{R}^{d},(s,a)\in\mathcal{S}\times\mathcal{A}$. Then there exists $c>0$ s.t. for any $\varepsilon>0,\delta\in(0,1)$ and for all $\theta\in\mathbb{R}^{d},$ we have $\|\hat{\nabla}_{n,m}J(\theta)-\nabla J(\theta)\|_{2}\leq\varepsilon$ with probability at least $1-\delta$, if $n\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(1/\delta)}{\varepsilon^{2}}$ and $m\geq\frac{(cHM_{\psi}M_{u}L_{w})^{2}\ln(2d/\delta)}{\varepsilon^{2}}\,.$
@@
We close this section by discussing the convergence of the policy parameter sequence produced by Algorithm 1. The CPT-RL objective is non-convex in the policy parameter due to non-convexity of both utility and probability weighting functions. While this lack of structure makes global optimality out of reach, we can still obtain a standard asymptotic convergence result towards the set of stationary points using the machinery of stochastic approximation (see e.g. Borkar (2008); Benaïm (2006)).